# Spin Bott Indices For Time-Reversal-Invariant Higher-Order Topological Superconductors

### v1

@@ Abstract
Discussion.— Higher-order topological phases have been classified into intrinsic type which hosts crystalline symmetry-protected bulk topology and extrinsic type whose topological phase transition can be tuned by closing boundary energy gap [4, 8]. Both the intrinsic and extrinsic HOTSCs can be realized by $\mathcal{H}$ in Eq. (11). In the case of $\lambda_{\text{R}}=0$ and $\Delta_{\text{A}}=-\Delta_{\text{B}}$, $\mathcal{H}$ describes an odd parity superconductor with an effective inversion symmetry in the bulk, where a $Z_{4}$ inversion symmetry indicator $\kappa$ can be defined as a bulk invariant to characterize the bulk topology [37, 63]. The intrinsic HOTSCs are realized in the three examples shown in Figs. 1(b),  2(a), and  2(e), where both the bulk and edge terminations preserve the inversion symmetry. Here the nontrivial bulk invariant $\kappa$ indicates the presence of MKP corner states for symmetry-preserving edge terminations but does not distinguish the different spatial patterns of MKP in the above three examples. In the case where the crystalline symmetry is broken in the bulk and/or by edge termination, extrinsic HOTSCs with time-reversal symmetry can also host robust MKP protected by (1) energy gaps of both bulk and boundaries and (2) symmetries of $P$ and $T$, as clearly demonstrated by results in Figs. 1 and 2. Our theory can characterize both intrinsic and extrinsic time-reversal invariant HOTSCs and distinguish different spatial patterns of MKP, while symmetry indicators such as $\kappa$ can not.
We further discuss the implications of our theory on other systems. In 1D superconductors, the spin Bott index in Eq. (6) can be defined by choosing the unitary matrix $M_{z}$ to be $e^{2\pi ixs_{z}\tau_{0}/L}$, which can be used to characterize the MKP of end states. Furthermore, our theory can also be generalized to superconductors in the BDI symmetry class, which host the chiral symmetry as a combination of $T$ and $P$ symmetries with $T^{2}=1$ and $P^{2}=1$. The BDI symmetry class can be realized in the superconducting systems with a magnetic group symmetry, such as $M_{x}T$ or $C_{2z}T$ symmetry [71, 72]. In this context, Majorana corner states protected by the chiral symmetry can be universally characterized by $N$ defined in Eq. (2). In superconductors belonging to the D symmetry class, where only the particle-hole symmetry $P$ is present, Majorana corner states are protected by this symmetry. The generalization of our theory to the D symmetry class is an interesting open question. In summary, we establish a new method for an exact characterization of HOTSCs using real-space invariants of spin Bott indices. This approach enhances the understanding of higher-order topological phases and opens up a range of applications.
@@
Acknowledgments.— X.-J. L. and F.W. are supported by Key Research and Development Program of Hubei Province (Grant No. 2022BAA017). J.-Z. L. and M.X. are supported by the National Key Research and Development Program of China (Grant No. 2022YFA1404900), the National Natural Science Foundation of China (Grant No. 12274332).
Supplemental Material for “Spin Bott Indices For Time-Reversal-Invariant Higher-Order Topological Superconductors”
This Supplemental Material includes the following seven sections: (1) equivalent definitions of the Bott index; (2) equivalent expressions of the spin Bott indices; (3) robustness of the spin Bott index; (4) edge theory; (5) Majoaran Kramers Pairs in the square- and dodecagonal-shaped systems; (6) absence of Majoaran Kramers Pairs; (7) $Z_{4}$ symmetry indicator.
I equivalent definitions of the Bott index
@@
$H=\begin{pmatrix}0&h\\ h^{\dagger}&0\end{pmatrix}.$ (S1)
@@
$N=\frac{1}{2\pi i}\text{Tr}\text{log}(mqm^{\dagger}q^{\dagger}),$ (S2)
@@
$MQM^{\dagger}Q$ $=$ $\begin{pmatrix}m&0\\ 0&m\end{pmatrix}\begin{pmatrix}0&q\\ q^{\dagger}&0\end{pmatrix}\begin{pmatrix}m^{\dagger}&0\\ 0&m^{\dagger}\end{pmatrix}\begin{pmatrix}0&q\\ q^{\dagger}&0\end{pmatrix}$ (S3)
@@
$N$ $=$ $\frac{1}{2\pi i}\text{Tr}\text{log}(mqm^{\dagger}q^{\dagger})$ (S4)
@@
$=$ $\frac{1}{4\pi i}\text{Tr}[C\text{log}(MQM^{\dagger}Q)].$ (S5)
In Eq. (S4), we have used the relation $\text{Trlog}(mqm^{\dagger}q^{\dagger})=-\text{Trlog}(qmq^{\dagger}m^{\dagger})=-\text{Trlog}(mq^{\dagger}m^{\dagger}q)$. $N$ can be also be defined as
$N$ $=$ $\frac{1}{2\pi i}\text{Tr}\text{log}(mqm^{\dagger}q^{\dagger})$ (S6)
@@
$=$ $\frac{1}{4\pi i}\text{Tr}\text{log}(\Xi).$ (S7)
where $\Xi=\frac{\mathbbm{1}+C}{2}{M}Q{M}^{\dagger}Q+\frac{\mathbbm{1}-C}{2}{M}^{\dagger}Q{M}Q$ and we have used the relation $\text{Trlog}(mqm^{\dagger}q^{\dagger})=\text{Trlog}(m^{\dagger}q^{\dagger}mq)$ in Eq. (S6). Thus, Eqs. (S2), (S5), and (S7) provide three equivalent definitions of the Bott index $N$.
II equivalent expressions of the spin Bott indices
@@
$\mathcal{N}$ $=$ $\frac{1}{4\pi i}\text{Tr}[\mathcal{C}\text{log}(MQM^{\dagger}Q)]$ (S8)
@@
$Q=\begin{pmatrix}Q_{+}&0\\ 0&Q_{-}\end{pmatrix}.$ (S9)
@@
$\mathcal{N}$ $=$ $\frac{1}{4\pi i}\text{Tr}[s_{z}C\text{log}(MQM^{\dagger}Q)$ (S10)
@@
where matrices $C_{\pm}$ and $M_{\pm}$ are spin resolved and have a dimension that is half of $C$ and $M$. In Eq. (S10), we have defined
$N_{\pm}=\frac{1}{4\pi i}\text{Tr}[C_{\pm}\text{log}(M_{\pm}Q_{\pm}M_{\pm}^{\dagger}Q_{\pm})].$ (S11)
@@
$\bar{\mathcal{N}}=\frac{1}{4\pi i}\text{Tr}[C\text{log}(M_{z}QM_{z}^{\dagger}Q)],$ (S12)
@@
$\bar{\mathcal{N}}$ $=$ $\frac{1}{4\pi i}\text{Tr}[C\text{log}(M_{z}QM_{z}^{\dagger}Q)$ (S13)
@@
$N_{-}$ $=$ $\frac{1}{4\pi i}\text{Tr}[C_{-}\text{log}(M_{-}Q_{-}M_{-}^{\dagger}Q_{-})]$ (S14)
@@
Besides Eq. (S8) and Eq. (S12), there is a third definition of spin Bott index by using the spin projection. We first construct the spin projection operator [73]
$P_{z}=Ps_{z}P,\quad P=(\mathbbm{1}-Q)/2,$ (S15)
@@
$P_{\pm}=\sum_{i}|\phi_{i}^{\pm}\rangle\langle\phi_{i}^{\pm}|,\quad\tilde{Q}_{\pm}=-P_{\pm}+CP_{\pm}C,$ (S16)
@@
$\tilde{\mathcal{N}}_{\pm}$ $=\frac{1}{4\pi i}\operatorname{Tr}[C\operatorname{log}(M\tilde{Q}_{\pm}M^{\dagger}\tilde{Q}_{\pm})]=\operatorname{Bott}(m,\tilde{q}_{\pm}),$ (S17)
$\tilde{\mathcal{N}}$ $=\tilde{\mathcal{N}}_{+}-\tilde{\mathcal{N}}_{-},$ (S18)
@@
When $[s_{z},H]=0$, the eigenvalues of $P_{z}$ consist of just two types of nonzero values, ±$\pm 1$. When projecting onto the spin up and down spaces, we have $\tilde{Q}_{\pm}=Q_{\pm}$, giving rise to $\tilde{\mathcal{N}}_{\pm}={\mathcal{N}}_{\pm}$ and $\tilde{\mathcal{N}}={\mathcal{N}}=\bar{\mathcal{N}}$. Thus, Eqs. (S8), (S12), and (S18) provide three equivalent expressions of the spin Bott index when $[s_{z},H]=0$.
III Robustness of the spin Bott index
In this section, we show that the Bott indices $\bar{\mathcal{N}}$ and $\tilde{\mathcal{N}}$, obtained through Eq. (S12) (Eq. (6) in the main text) and Eq. (S18), respectively, are robust, and $\bar{\mathcal{N}}=\tilde{\mathcal{N}}$, unless the energy gap of the system closes or the gap of $P_{z}$ closes.
III.1 Notations
@@
$\parallel A+B\parallel$ $\leq\parallel A\parallel+\parallel B\parallel$ (S19)
$\parallel AB\parallel$ $\leq\parallel A\parallel\parallel B\parallel.$ (S20)
@@
III.2 Mathematical derivation
First, we consider the spin Bott index in Eq. (S12). We have
$\bar{\mathcal{N}}=\frac{1}{4\pi i}\text{Tr}[C\text{log}(M_{z}QM_{z}^{\dagger}Q)]=\operatorname{Bott}(m_{z},q),$ (S21)
@@ III.2 Mathematical derivation
$\operatorname{Bott}\left(V(s),W(s)\right)=\operatorname{Bott}(V,W).$ (S23)
@@
$\parallel[m_{z},q]\parallel$ $\leq\parallel[c_{z}e^{-2\pi if(\bm{r})},q]\parallel+\parallel[(1-c_{z})e^{2\pi if(\bm{r})},q]\parallel$ (S24)
@@
$H=\begin{pmatrix}0&h\\ h^{\dagger}&0\end{pmatrix},$ (S25)
@@
$\parallel[m,q]\parallel$ $=\parallel mqm^{\dagger}q^{\dagger}-1\parallel$ (S28)
@@
$U_{A,\text{corner}}^{\dagger}mU_{A,\text{corner}}U_{B,\text{corner}}^{\dagger}m^{\dagger}U_{B,\text{corner}}=\mathbbm{1}$ (S29)
@@
$\parallel U_{A}^{\dagger}mU_{A}U_{B}^{\dagger}m^{\dagger}U_{B}-1\parallel$ $=\parallel\bigoplus_{\beta\in\{\text{bulk},\text{edge},\text{corner}\}}U_{A,\beta}^{\dagger}mU_{A,\beta}U_{B,\beta}^{\dagger}m^{\dagger}U_{B,\beta}-1\parallel$ (S30)
@@
$\operatorname{sup}_{\beta\in\{\text{bulk},\text{edge}\}}(\parallel U_{A,\beta}^{\dagger}mU_{A,\beta}U_{B,\beta}^{\dagger}m^{\dagger}U_{B,\beta}-1\parallel)\leq\mathcal{O}\left(\frac{R}{L}\frac{\parallel H\parallel}{\Delta E}\right).$ (S31)
@@
$\parallel[m,q]\parallel\leq\mathcal{O}\left(\frac{R}{L}\frac{\parallel H\parallel}{\Delta E}\right).$ (S32)
@@
$\parallel[m_{z},q]\parallel$ $\leq 2\parallel[c_{z},q]\parallel+\mathcal{O}\left(\frac{R}{L}\frac{\parallel H\parallel}{\Delta E}\right)$ (S33)
@@
$P=\left(\begin{matrix}\frac{1}{2}&-\frac{q}{2}\\ -\frac{q^{\dagger}}{2}&\frac{1}{2}\\ \end{matrix}\right).$ (S34)
@@
$P=\begin{pmatrix}1&0\\ 0&0\end{pmatrix},$ (S35)
@@
$s_{z}=\begin{pmatrix}A&D\\ D^{\dagger}&B\end{pmatrix},$ (S36)
@@
$[s_{z},P]=\begin{pmatrix}0&-D\\ D^{\dagger}&0\end{pmatrix},$ (S38)
@@
$P_{z}=Ps_{z}P=\begin{pmatrix}A&0\\ 0&0\end{pmatrix}.$ (S39)
@@
$[s_{z},P]^{\dagger}[s_{z},P]=\begin{pmatrix}DD^{\dagger}&0\\ 0&DD^{\dagger}\end{pmatrix}.$ (S40)
@@
$\sigma_{max}([s_{z},P]^{\dagger}[s_{z},P])=\sigma_{max}(DD^{\dagger}).$ (S41)
@@
$\langle\psi_{max}|A^{2}+DD^{\dagger}|\psi_{max}\rangle=1=\langle\psi_{max}|A^{2}|\psi_{max}\rangle+\sigma_{max}.$ (S42)
@@
$\sigma_{max}=1-\langle\psi_{max}|A^{2}|\psi_{max}\rangle\leq 1,$ (S43)
@@
$\parallel[m_{z},q]\parallel<2.$ (S44)
According to Theorem S1, we have demonstrated that the Bott index $\bar{\mathcal{N}}$ is robust with the existence of both a gap of $P_{z}$ at zero and $\Delta E$.
Next, we consider the spin Bott index defined by Eq. (S18).
Utilizing Theorem S2, we can prove the following theorem:
Theorem S3.
@@
$\left[A,(P_{z}-z)^{-1}\right]$ $=(P_{z}-z)^{-1}\left[(P_{z}-z),A\right](P_{z}-z)^{-1}$ (S48)
@@
$\parallel[m,P_{-}]\parallel\leq\frac{1}{2\pi}\parallel[m,P_{z}]\parallel\oint_{\Gamma}\parallel\left(P_{z}-z\right)^{-1}\parallel^{2}|dz|,$ (S49)
@@
$\parallel[m,P_{z}]\parallel$ $=\parallel[m,P]s_{z}P+Ps_{z}[m,P]\parallel$ (S51)
@@
where we use two inequalities of matrix norm Eqs. (S19, S20), and in the last step we use Theorem S2.
@@
$\parallel[m,P_{-}]\parallel\leq\mathcal{O}\left(\frac{R}{L}\frac{\parallel H\parallel}{\Delta E\Delta E_{z}}\right),$ (S52)
@@
Applying this theorem to Eq. (S16), we have
$\parallel[M,\tilde{Q}_{\pm}]\parallel\leq\mathcal{O}\left(\frac{R}{L}\frac{\parallel H\parallel}{\Delta E\Delta E_{z}}\right),$ (S53)
@@
$H=H_{\text{cons}}+H_{\text{brok}},$ (S54)
@@
$H(s)=H_{\text{cons}}+sH_{\text{brok}},\quad s\in[0,1].$ (S55)
@@
when $L\to\infty$. $q(s)$ and $\tilde{q}_{\pm}(s)$ are obtained by replacing $H$ with $H(s)$. Applying Theorem S1 to $\bar{N}(s)=\operatorname{Bott}(m_{z},q(s))$ and $\tilde{N}_{\pm}(s)=\operatorname{Bott}(m,\tilde{q}_{\pm}(s))$, we have
$\bar{N}(1)=\bar{N}(0),$ (S57)
@@
$\bar{N}(1)=\bar{N}(0)=\tilde{N}(0)=\tilde{N}_{+}(0)-\tilde{N}_{-}(0)=\tilde{N}_{+}(1)-\tilde{N}_{-}(1)=\tilde{N}(1).$ (S58)
@@
Figure S1: Energy spectra for a ribbon with armchair edge in (a) and (b), zigzag edge in (c) and (d), and bearded edge in (e) and (f). The common parameters are taken as $t=1,\lambda_{\text{so}}=0.1,\Delta_{\text{A}}=0.2$, and $\mu=\lambda_{v}=\lambda_{\text{R}}=0$.
IV Edge theory

### v3

@@ VI Discussion and summary
Higher-order topological phases have been classified into intrinsic type which hosts crystalline symmetry-protected bulk topology and extrinsic type whose topological phase transition can be tuned by closing boundary energy gap [4, 8]. Both the intrinsic and extrinsic HOTSCs can be realized by $\mathcal{H}$ in Eq. (11). In the case of $\lambda_{\text{R}}=0$ and $\Delta_{\text{A}}=-\Delta_{\text{B}}$, $\mathcal{H}$ describes an odd parity superconductor with an effective inversion symmetry in the bulk, where a $Z_{4}$ inversion symmetry indicator [40] $\kappa$ can be defined as a bulk invariant to characterize the bulk topology (see Appendix G for details). The intrinsic HOTSCs are realized in the three examples shown in Figs. 1(b),  2(a), and  2(e), where both the bulk and edge terminations preserve the inversion symmetry. Here the nontrivial bulk invariant $\kappa$ indicates the presence of MKP corner states for symmetry-preserving edge terminations but does not distinguish the different spatial patterns of MKP in the above three examples. In the case where the crystalline symmetry is broken in the bulk and/or by edge termination, extrinsic HOTSCs with time-reversal symmetry can also host robust MKP protected by (1) energy gaps of both bulk and boundaries and (2) symmetries of $P$ and $T$, as clearly demonstrated by results in Figs. 1 and 2. Our theory can characterize both intrinsic and extrinsic time-reversal invariant HOTSCs and distinguish different spatial patterns of MKP, while symmetry indicators such as $\kappa$ cannot.
In superconductors belonging to the D symmetry class, where only the particle-hole symmetry $P$ is present, Majorana corner states are protected by this symmetry. This scenario can be realized, for example, by further adding a Rashba term to Eq. (13), where a small $\lambda_{\text{R}}$ cannot remove the Majorana corner states but can change the symmetry classes. The generalization of our theory to the D symmetry class is an open question.
In summary, we establish a real-space method for characterization of HOTSCs in both DIII and BDI symmetry classes using invariants of Bott indices. We apply our theory to characterize various patterns of Majorana corner states in a representative model with different shapes. Our study introduces Bott indices for characterizing higher-order topological superconductors, which advances the understanding of their properties and opens up a range of applications.
VII Acknowledgments
@@
X.-J. L. and F.W. are supported by the National Key Research and Development Program of China (Grant No. 2022YFA1402400) and the National Natural Science Foundation of China (Grant No. 12274333). J.-Z. L. and M.X. are supported by the National Key Research and Development Program of China (Grant No. 2022YFA1404900), the National Natural Science Foundation of China (Grant No. 12274332).
Appendix A equivalent definitions of the Bott index
@@
$H=\begin{pmatrix}0&h\\ h^{\dagger}&0\end{pmatrix}.$ (14)
@@
$N=\frac{1}{2\pi i}\text{Tr}\text{log}(mqm^{\dagger}q^{\dagger}),$ (15)
@@
$MQM^{\dagger}Q$ $=$ $\begin{pmatrix}m&0\\ 0&m\end{pmatrix}\begin{pmatrix}0&q\\ q^{\dagger}&0\end{pmatrix}\begin{pmatrix}m^{\dagger}&0\\ 0&m^{\dagger}\end{pmatrix}\begin{pmatrix}0&q\\ q^{\dagger}&0\end{pmatrix}$ (16)
@@
$N$ $=$ $\frac{1}{2\pi i}\text{Tr}\text{log}(mqm^{\dagger}q^{\dagger})$ (17)
@@
$=$ $\frac{1}{4\pi i}\text{Tr}[C\text{log}(MQM^{\dagger}Q)].$ (18)
In Eq. (17), we have used the relation $\text{Trlog}(mqm^{\dagger}q^{\dagger})=-\text{Trlog}(qmq^{\dagger}m^{\dagger})=-\text{Trlog}(mq^{\dagger}m^{\dagger}q)$. $N$ can be also be defined as
$N$ $=$ $\frac{1}{2\pi i}\text{Tr}\text{log}(mqm^{\dagger}q^{\dagger})$ (19)
@@
$=$ $\frac{1}{4\pi i}\text{Tr}\text{log}(\Xi).$ (20)
where $\Xi=\frac{\mathbbm{1}+C}{2}{M}Q{M}^{\dagger}Q+\frac{\mathbbm{1}-C}{2}{M}^{\dagger}Q{M}Q$ and we have used the relation $\text{Trlog}(mqm^{\dagger}q^{\dagger})=\text{Trlog}(m^{\dagger}q^{\dagger}mq)$ in Eq. (19). Thus, Eqs. (15), (18), and (20) provide three equivalent definitions of the Bott index $N$.
Appendix B equivalent expressions of the spin Bott indices
@@
$\mathcal{N}$ $=$ $\frac{1}{4\pi i}\text{Tr}[\mathcal{C}\text{log}(MQM^{\dagger}Q)]$ (21)
@@
$Q=\begin{pmatrix}Q_{+}&0\\ 0&Q_{-}\end{pmatrix}.$ (22)
@@
$\mathcal{N}$ $=$ $\frac{1}{4\pi i}\text{Tr}[s_{z}C\text{log}(MQM^{\dagger}Q)$ (23)
@@
where matrices $C_{\pm}$ and $M_{\pm}$ are spin resolved and have a dimension that is half of $C$ and $M$. In Eq. (23), we have defined
$N_{\pm}=\frac{1}{4\pi i}\text{Tr}[C_{\pm}\text{log}(M_{\pm}Q_{\pm}M_{\pm}^{\dagger}Q_{\pm})].$ (24)
@@
$\bar{\mathcal{N}}=\frac{1}{4\pi i}\text{Tr}[C\text{log}(M_{z}QM_{z}^{\dagger}Q)],$ (25)
@@
$\bar{\mathcal{N}}$ $=$ $\frac{1}{4\pi i}\text{Tr}[C\text{log}(M_{z}QM_{z}^{\dagger}Q)$ (26)
@@
$N_{-}$ $=$ $\frac{1}{4\pi i}\text{Tr}[C_{-}\text{log}(M_{-}Q_{-}M_{-}^{\dagger}Q_{-})]$ (27)
@@
Besides Eq. (21) and Eq. (25), there is a third definition of spin Bott index by using the spin projection. We first construct the spin projection operator [89]
$P_{z}=Ps_{z}P,\quad P=(\mathbbm{1}-Q)/2,$ (28)
@@
$P_{\pm}=\sum_{i}|\phi_{i}^{\pm}\rangle\langle\phi_{i}^{\pm}|,\quad\tilde{Q}_{\pm}=-P_{\pm}+CP_{\pm}C,$ (29)
@@
$\tilde{\mathcal{N}}_{\pm}$ $=\frac{1}{4\pi i}\operatorname{Tr}[C\operatorname{log}(M\tilde{Q}_{\pm}M^{\dagger}\tilde{Q}_{\pm})]=\operatorname{Bott}(m,\tilde{q}_{\pm}),$ (30)
$\tilde{\mathcal{N}}$ $=\tilde{\mathcal{N}}_{+}-\tilde{\mathcal{N}}_{-},$
@@
When $[s_{z},H]=0$, the eigenvalues of $P_{z}$ consist of just two types of nonzero values, ±$\pm 1$. When projecting onto the spin up and down spaces, we have $\tilde{Q}_{\pm}=Q_{\pm}$, giving rise to $\tilde{\mathcal{N}}_{\pm}={\mathcal{N}}_{\pm}$ and $\tilde{\mathcal{N}}={\mathcal{N}}=\bar{\mathcal{N}}$. Thus, Eqs. (21), (25), and (30) provide three equivalent expressions of the spin Bott index when $[s_{z},H]=0$.
Appendix C Alternative choice of polynomials $f_{6}^{(i)}$
The choice of polynomials $f_{p}^{(i)}$ is not unique. For example, when $p=6$, besides the given expression in Eq. (9), $f_{6}^{(i)}$ can be alternatively chosen as
$f_{6}^{(1)}(r)=\frac{\left(x^{3}-\frac{8x^{2}y}{\sqrt{3}}-3xy^{2}+\frac{8y^{3}}{3\sqrt{3}}\right)}{2L^{3}},$
$f_{6}^{(2)}(r)=\frac{\left(x^{3}-\frac{8x^{2}y}{\sqrt{3}}-\frac{xy^{2}}{3}\right)}{2L^{3}},$
$f_{6}^{(3)}(r)=\frac{\left(x^{3}-\frac{8x^{2}y}{\sqrt{3}}-\frac{xy^{2}}{3}+\frac{16y^{3}}{3\sqrt{3}}\right)}{2L^{3}},$
$f_{6}^{(4)}(r)=\frac{\left(x^{2}-\frac{5y^{2}}{3}\right)}{2L^{2}},$
$f_{6}^{(5)}(r)=\frac{\left(x^{2}-\frac{4xy}{\sqrt{3}}-\frac{y^{2}}{3}\right)}{2L^{2}},$ (31)
The above polynomials give rise to $\text{det}(M)\neq 0$.
Appendix D Robustness of the spin Bott index
In this section, we show that the Bott indices $\bar{\mathcal{N}}$ and $\tilde{\mathcal{N}}$, obtained through Eq. (25) (Eq. (6) in the main text) and Eq. (30), respectively, are robust, and $\bar{\mathcal{N}}=\tilde{\mathcal{N}}$, unless the energy gap of the system closes or the gap of $P_{z}$ closes.
D.1 Notations
@@
$\parallel A+B\parallel$ $\leq\parallel A\parallel+\parallel B\parallel$ (32)
$\parallel AB\parallel$ $\leq\parallel A\parallel\parallel B\parallel.$ (33)
@@
D.2 Mathematical derivation
First, we consider the spin Bott index in Eq. (25). We have
$\bar{\mathcal{N}}=\frac{1}{4\pi i}\text{Tr}[C\text{log}(M_{z}QM_{z}^{\dagger}Q)]=\operatorname{Bott}(m_{z},q),$ (34)
@@ D.2 Mathematical derivation
$\operatorname{Bott}\left(V(s),W(s)\right)=\operatorname{Bott}(V,W).$ (36)
@@
$\parallel[m_{z},q]\parallel$ $\leq\parallel[c_{z}e^{-2\pi if(\bm{r})},q]\parallel+\parallel[(1-c_{z})e^{2\pi if(\bm{r})},q]\parallel$ (37)
@@
$H=\begin{pmatrix}0&h\\ h^{\dagger}&0\end{pmatrix},$ (38)
@@
$\parallel[m,q]\parallel$ $=\parallel mqm^{\dagger}q^{\dagger}-1\parallel$ (41)
@@
$U_{A,\text{corner}}^{\dagger}mU_{A,\text{corner}}U_{B,\text{corner}}^{\dagger}m^{\dagger}U_{B,\text{corner}}=\mathbbm{1}$ (42)
@@
$\parallel U_{A}^{\dagger}mU_{A}U_{B}^{\dagger}m^{\dagger}U_{B}-1\parallel$ $=\parallel\bigoplus_{\beta\in\{\text{bulk},\text{edge},\text{corner}\}}U_{A,\beta}^{\dagger}mU_{A,\beta}U_{B,\beta}^{\dagger}m^{\dagger}U_{B,\beta}-1\parallel$ (43)
@@
$\operatorname{sup}_{\beta\in\{\text{bulk},\text{edge}\}}(\parallel U_{A,\beta}^{\dagger}mU_{A,\beta}U_{B,\beta}^{\dagger}m^{\dagger}U_{B,\beta}-1\parallel)\leq\mathcal{O}\left(\frac{R}{L}\frac{\parallel H\parallel}{\Delta E}\right).$ (44)
@@
$\parallel[m,q]\parallel\leq\mathcal{O}\left(\frac{R}{L}\frac{\parallel H\parallel}{\Delta E}\right).$ (45)
@@
$\parallel[m_{z},q]\parallel$ $\leq 2\parallel[c_{z},q]\parallel+\mathcal{O}\left(\frac{R}{L}\frac{\parallel H\parallel}{\Delta E}\right)$ (46)
@@
$P=\left(\begin{matrix}\frac{1}{2}&-\frac{q}{2}\\ -\frac{q^{\dagger}}{2}&\frac{1}{2}\\ \end{matrix}\right).$ (47)
@@
$P=\begin{pmatrix}1&0\\ 0&0\end{pmatrix},$ (48)
@@
$s_{z}=\begin{pmatrix}A&D\\ D^{\dagger}&B\end{pmatrix},$ (49)
@@
$[s_{z},P]=\begin{pmatrix}0&-D\\ D^{\dagger}&0\end{pmatrix},$ (51)
@@
$P_{z}=Ps_{z}P=\begin{pmatrix}A&0\\ 0&0\end{pmatrix}.$ (52)
@@
$[s_{z},P]^{\dagger}[s_{z},P]=\begin{pmatrix}DD^{\dagger}&0\\ 0&DD^{\dagger}\end{pmatrix}.$ (53)
@@
$\sigma_{max}([s_{z},P]^{\dagger}[s_{z},P])=\sigma_{max}(DD^{\dagger}).$ (54)
@@
$\langle\psi_{max}|A^{2}+DD^{\dagger}|\psi_{max}\rangle=1=\langle\psi_{max}|A^{2}|\psi_{max}\rangle+\sigma_{max}.$ (55)
@@
$\sigma_{max}=1-\langle\psi_{max}|A^{2}|\psi_{max}\rangle\leq 1,$ (56)
@@
$\parallel[m_{z},q]\parallel<2.$ (57)
According to Theorem 1, we have demonstrated that the Bott index $\bar{\mathcal{N}}$ is robust with the existence of both a gap of $P_{z}$ at zero and $\Delta E$.
Figure 5: Energy spectra for a ribbon with armchair edge in (a) and (b), zigzag edge in (c) and (d), and bearded edge in (e) and (f). The common parameters are taken as $t=1,\lambda_{\text{so}}=0.1,\Delta_{\text{A}}=0.2$, and $\mu=\lambda_{v}=\lambda_{\text{R}}=0$.
Next, we consider the spin Bott index defined by Eq. (30).
Utilizing Theorem 2, we can prove the following theorem:
Theorem 3.
@@
$\left[A,(P_{z}-z)^{-1}\right]$ $=(P_{z}-z)^{-1}\left[(P_{z}-z),A\right](P_{z}-z)^{-1}$ (61)
@@
$\parallel[m,P_{-}]\parallel\leq\frac{1}{2\pi}\parallel[m,P_{z}]\parallel\oint_{\Gamma}\parallel\left(P_{z}-z\right)^{-1}\parallel^{2}|dz|,$ (62)
@@
$\parallel[m,P_{z}]\parallel$ $=\parallel[m,P]s_{z}P+Ps_{z}[m,P]\parallel$ (64)
@@
where we use two inequalities of matrix norm Eqs. (32, 33), and in the last step we use Theorem 2.
@@
$\parallel[m,P_{-}]\parallel\leq\mathcal{O}\left(\frac{R}{L}\frac{\parallel H\parallel}{\Delta E\Delta E_{z}}\right),$ (65)
@@
Applying this theorem to Eq. (29), we have
$\parallel[M,\tilde{Q}_{\pm}]\parallel\leq\mathcal{O}\left(\frac{R}{L}\frac{\parallel H\parallel}{\Delta E\Delta E_{z}}\right),$ (66)
@@
$H=H_{\text{cons}}+H_{\text{brok}},$ (67)
@@
$H(s)=H_{\text{cons}}+sH_{\text{brok}},\quad s\in[0,1].$ (68)
@@
when $L\to\infty$. $q(s)$ and $\tilde{q}_{\pm}(s)$ are obtained by replacing $H$ with $H(s)$. Applying Theorem 1 to $\bar{N}(s)=\operatorname{Bott}(m_{z},q(s))$ and $\tilde{N}_{\pm}(s)=\operatorname{Bott}(m,\tilde{q}_{\pm}(s))$, we have
$\bar{N}(1)=\bar{N}(0),$ (70)
@@
$\bar{N}(1)=\bar{N}(0)$ $=\tilde{N}(0)$ (71)
$=\tilde{N}_{+}(0)-\tilde{N}_{-}(0)$
$=\tilde{N}_{+}(1)-\tilde{N}_{-}(1)$
$=\tilde{N}(1).$
@@
Figure 6: Different patterns of MKP on the square-shaped systems. The circled number in (a) at corners labels the order of corners. The letters $\text{A}_{\text{AB}}$, $\text{Z}_{\alpha}$, $\text{B}_{\alpha}$ label the armchair, zigzag, and bearded edges, respectively, with $\alpha\in$A, B}. The insets plot the energies close to zero. The vectors $\bm{\chi}_{4}$ and $\bm{\mathcal{N}}_{4}$ are used to describe and characterize the different patterns of MKP. The model parameters are taken as $t=1,\lambda_{so}=0.2,\Delta_{\text{A}}=0.5,\Delta_{\text{B}}=-0.25$.
Appendix E Edge theory